ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

17. Control Theory and Optimization

1991 MS Classification: 93, 49, 34, 47 Obukhovskii, Valeri, Voronezh Pedagogical University, Russia On some methods in a periodic problem for a functional differential inclusion in a Banach space We consider a problem of existence of periodic solutions for a system governed by a semilinear functional differential inclusion in a Banach space. The properties of the translation multioperator along the trajectories of the inclusion are studied. On this base application of a new asymptotic fixed point theorem yields the existence of a periodic solution for a dissipative inclusion. The presence of a global compact attractor for the set of solutions is also demonstrated. Part of these results is joint work with P. Zecca. Reference: V. Obukhovskii and P. Zecca, On some properties of dissipative functional differential inclusions in a Banach space, to appear in Topological Methods of Nonlinear Analysis.